Question 5: For the Givens-Rotator Q the determinant. (8) ER. (see page 194 of the text book) show that det(Q), 8)) 1 by ex194 THE LEAST SQUARES PROBLEM We now turn our attention to n x n matrices. A plane rotator is a matrix of the form - S Q= (3.
Answer #1

Illici 5. Qij (0,89 ИХИ Cocoso, so sind. det (Qij (0,8)) = 11- 1x1xl So..-5 1-1 Clio: 01 - 0 So.-... n- 101-01 Tool. O . 1+1 enem = c.cas. (-5) = c² + s2 = cos o + sin 0= 1. i + 1 odd, t